# Algebraic flat connections and o-minimality

### v1

@@ 1. Introduction
Let $X$ be a smooth complex algebraic variety. Let $\mathcal{E}$ be an algebraic vector bundle on $X$ and $\nabla\colon\mathcal{E}\to\Omega^{1}_{X}\otimes_{\mathcal{O}_{X}}\mathcal{E}$ be a flat algebraic connection. We denote the analytic spaces underlying $X$ and $\mathcal{E}$ by the same symbol and we endow them with their canonical $\mathbb{R}_{\rm an,exp}$-definable structure, see [BBT23]. Roughly speaking, a function on $X$ is definable in this structure if it can be expressed in terms of the usual logical operators (including quantifiers), algebraic functions, real analytic functions on compact domains like $[0,1]^{n}$ and the real exponential function. A fundamental property of this structure is its o-minimality, which essentially means that definable functions don’t exhibit oscillatory behavior.
@@
    The ‘if’ part is due to Bakker–Mullane [BM23] and the ‘only if’ part refines their Example 3.3.
@@
    One can easily show that the property of definable flat sections holds for all definable open subsets if and only if it holds for a single finite covering by simply connected open definable subsets (see the proof of Lemma 4.1). In addition, all flat sections on a given definable $U$ are definable if and only it is true on a basis of such.
@@
      By requiring that for all non-singular complex curves $C$ and all morphisms $\psi\colon C\to X$ the flat connection $\psi^{*}(\mathcal{E},\nabla)$ has monodromy eigenvalues of absolute value one at each point of the normal compactification $\overline{C}$ of $C$.
@@
The algebraic de Rham cohomology together with the Gauss-Manin connection $(\mathcal{H}_{\rm dR}^{j}(Y/X),\nabla)$ of a smooth projective family $Y\to X$ of complex algebraic varieties has definable flat sections.
@@
In the non-linear setting of non-abelian Hodge theory it is not clear to us whether the answer is always positive or whether there is any direct connection to Simpson’s observations about the non-abelian Gauss-Manin connection in the spirit of Theorem 1.1.
Acknowledgment. It is a pleasure to acknowledge discussions around the topic of this note with Benjamin Bakker, Philip Engel, Bruno Klingler and Salim Tayou.
@@ 2. Reminder on linear meromorphic differential equations
on the open disc $\Delta(\rho)=\{z\in{\mathbb{C}},|z|<\rho\}$. Here $A$ is an $r\times r$-matrix with entries holomorphic functions on the punctured disc $\Delta(\rho)^{\times}=\Delta(\rho)\setminus\{0\}$ which are meromorphic along 00. We assume that $A$ is overconvergent, i.e. that all its entries extend to holomorphic functions on $\Delta(\tilde{\rho})^{\times}$ for some $\tilde{\rho}>\rho$.
Recall that a gauge transformation by $P$ changes $A$ to the coefficient matrix $P^{-1}AP-zP^{-1}P^{\prime}$ in (2.1). A pullback along the finite covering $z\mapsto z^{m}$ for $m\in\mathbb{Z}_{\geq 1}$ replaces $A(z)$ by $mA(z^{m})$. We say that $A$ is in split normal form if there exist $Q_{1}(z),\ldots,Q_{\ell}(z)\in\frac{1}{z}{\mathbb{C}}[\frac{1}{z}]$ and constant square matrices $B_{1},\ldots,B_{\ell}$ of size $r_{1},\ldots,r_{\ell}$ with $\sum_{i=1}^{\ell}r_{i}=r$, such that $A$ is block diagonal with blocks $Q_{1}{\rm Id}_{r_{1}}+B_{1},\ldots,Q_{\ell}{\rm Id}_{r_{\ell}}+B_{\ell}$. Note that for such a split normal form $A$ a fundamental matrix solution on a sector has the shape
@@
  • $f$ is the restriction to $D_{g}(\rho)$ of a real analytic function defined on $D_{d}(\tilde{\rho})$ for some $\tilde{\rho}>\rho$;
@@
    for all $n\geq 0$ and all $z$ in this sectorial domain;
  • the Taylor coefficients $\lim_{z\to 0}f_{j}^{(n)}(z)$ exist for the above $f_{j}$ for all $n\geq 0$.
@@
  • The condition on the slopes in Theorem 2.3 is satisfied after pullback along a finite covering $z\mapsto z^{m}$ for $m$ large.
@@
is a fundamental matrix solution of the differential equation (2.1) on $D_{d}$.
@@
Note here there is a change of notation as compared to (2.2), $\ell$ becomes $L$. This is because we request the $B_{i}$ to be Jordan blocks, which is achieved after a constant gauge transformation starting from the $B_{i},Q_{i}(z)$ as in (2.2). So in fact $L\geq\ell$.
@@ 3. Reminder on o-minimal structures
Recall that an o-minimal structure $(\mathcal{A}_{n})_{n}$ consists of a sets of subsets $A_{n}$ of $\mathbb{R}^{n}$ which satisfy certain properties, in particular they comprise the semi-algebraic sets and $\mathcal{A}_{1}$ consists of the finite unions of intervals. In this note if not explicitly mentioned otherwise we use the o-minimal structure $\mathbb{R}_{\rm an,exp}$ which is the structure generated by the semi-algebraic sets, the graphs of real analytic functions on $[0,1]^{n}$ and the graph of $\exp\colon\mathbb{R}\to\mathbb{R}$. The key property of o-minimal structures that we use is that a real analytic definable function $f\colon(0,1)\to\mathbb{R}$ which is not identically zero has only a finite number of zeros. A basic reference for o-minimal structures is [vD98].
@@
By an overconvergent flat bundle $(\mathcal{E},\nabla)$ on $U(\rho)=(\Delta(\rho)^{\times})^{n}\times\Delta(\rho)^{d-n}$ meromorphic along $D(\rho)=\Delta(\rho)^{d}\setminus U(\rho)$ we mean the restriction to $U(\rho)$ of a flat bundle on $U(\tilde{\rho})$ for some $\tilde{\rho}>\rho$ with a meromorphic structure along $D(\tilde{\rho})$ in the sense of [Del70, II.2.13], i.e. it is provided with a prolongation to a coherent sheaf of $\mathcal{M}_{U(\tilde{\rho}),D(\tilde{\rho})}$-modules with flat connection, where $\mathcal{M}_{U(\tilde{\rho}),D(\tilde{\rho})}$ is the sheaf of rings of holomorphic functions on $U(\tilde{\rho})$ which are meromorphic along $D(\tilde{\rho})$. Such a $\mathcal{E}$ has a canonical definable structure.
@@
The elementary technical observation about oscillating functions that we need in the proof of our main theorem is the following. For a complex number $z$ we write $z={\rm Re}(z)+i{\rm Im}(z),\ {\rm Re}(z),{\rm Im}(z)\in{\mathbb{R}}$.
@@
and grows more rapidly than $|{\rm log}(z)|$. Thus
@@ 4. Reduction to sectors
We only have to show $\Leftarrow$”. Let $U\subset X$ be an arbitrary definable open subset and $\sigma\colon U\to\mathcal{E}|_{U}$ a flat section. Let $\overline{U}\subset\overline{X}$ be the closure of $U$. Then there are finitely many charts $\phi$ as above covering $\overline{U}$ and finitely many polysectors $S$ in these charts which cover $U$. By definability of $U$, for any such polysector $S$, the number of connected components of $U\cap S$ is finite [vD98, Sec. 2.2]. Let $S^{\circ}$ be such a connected component. We can extend the flat section $\sigma|_{S^{\circ}}$ to a flat section of $\mathcal{E}|_{S}$ which is definable by assumption. So also $\sigma|_{S^{\circ}}$ is definable and therefore so is $\sigma_{U\cap S}$. ∎
@@
$\Leftarrow$” (Bakker–Mullane) By assumption the eigenvalues of $M_{1},\ldots,M_{n}$ are unitary. Let $B_{1},\ldots,B_{n}$ be commuting complex matrices with real eigenvalues and $M_{j}=e^{2\pi iB_{j}}$. Then
@@
is meromorphic along $D$, i.e. $f_{j}\equiv 0$ for $j\ll 0$. For this define $E_{j}=\{f_{j}=0\}\subset\Delta(\rho)^{d-1}$ if $f_{i}\not\equiv 0$ and $E_{j}=\varnothing$ else. This is a meager subset of $\Delta(\rho)^{d-1}$. So the countable union $\cup_{j}E_{j}$ is a meagre subset of $\Delta(\rho)^{d-1}$ as well. Thus there exists $(\tilde{z}_{2},\ldots,\tilde{z}_{d})\in\Delta(\rho)^{d-1}\setminus E$. Then if $\mathcal{\tilde{Y}}(z)$ is not meromorphic along $D$, the one variable function $z_{1}\mapsto\mathcal{\tilde{Y}}(z_{1},\tilde{z}_{2},\ldots,\tilde{z}_{d})$ is not meromorphic either. This reduces us to the case $d=1$ for checking regularity. The monodromy $M_{1}$ is also calculated by restriction to these one-dimensional discs. So the implication follows from Proposition 5.1
@@ 5. Local one variable case
The coherent sheaf of $\mathcal{M}_{(\Delta(\rho)^{\times},0)}$-modules $\mathcal{E}$ on $\Delta(\rho)$ is then locally free around 00, so without loss of generality $\mathcal{E}$ is the free sheaf ($\mathcal{M}^{r}_{(\Delta(\rho),0)}$. Then $(\mathcal{E},\nabla)$ corresponds to an overconvergent meromorphic differential equation (2.1), see [Del70, Sec. I.3].
Assume Proposition 5.1 is false, and say the connection is irregular singular. We can take the pullback along $z\mapsto z^{m}$ for $m$ large and perform a meromorphic gauge transformation so that we are in the situation of Corollary 2.5. Assume without loss of generality that the $Q_{1}(z)$ in this corollary is non-zero, say
@@
we obtain the the $\mathbb{R}_{\rm sum,exp}$-definable function

### v2

@@ 1. Introduction
Let $X$ be a smooth complex algebraic variety. Let $\mathcal{E}$ be an algebraic vector bundle on $X$ and $\nabla\colon\mathcal{E}\to\Omega^{1}_{X}\otimes_{\mathcal{O}_{X}}\mathcal{E}$ be a flat algebraic connection. We denote the analytic spaces underlying $X$ and $\mathcal{E}$ by the same symbol and we endow them with their canonical $\mathbb{R}_{\rm an,exp}$-definable structure, see [BBT23, Sec. 2 and 3]. Roughly speaking, a function on $X$ is definable in this structure if it can be expressed in terms of the usual logical operators (including quantifiers), algebraic functions, real analytic functions on compact domains like $[0,1]^{n}$ and the real exponential function. A fundamental property of this structure is its o-minimality, which essentially means that definable functions don’t exhibit oscillatory behavior.
@@
    The ‘if’ part is due to Bakker–Mullane [BM23, Thm. 1.2] and the ‘only if’ part refines their Example 3.3.
@@
    One can easily show that the property of definable flat sections holds for all definable open subsets if and only if it holds for a single finite covering by simply connected open definable subsets (see the proof of Lemma 4.1). In addition, all flat sections on a given definable $U$ are definable if and only this property is true on a basis of such.
@@
      By requiring that for all non-singular complex curves $C$ and all morphisms $\psi\colon C\to X$ the flat connection $\psi^{*}(\mathcal{E},\nabla)$ has monodromy eigenvalues of absolute value one at each point of the smooth compactification $\overline{C}$ of $C$.
@@
()($\star$)( ⋆ )    The algebraic de Rham cohomology together with the Gauss-Manin connection $(\mathcal{H}_{\rm dR}^{j}(Y/X),\nabla)$ of a smooth projective family $Y\to X$ of complex algebraic varieties has definable flat sections.
@@
In the non-linear setting of non-abelian Hodge theory it is not clear to us whether all flat sections are definable or whether there is any direct connection to Simpson’s regularity result for the non-abelian Gauss-Manin connection [Sim97, Sec. 8] in the spirit of Theorem 1.1.
Acknowledgments. It is a pleasure to acknowledge discussions around the topic of this note with Benjamin Bakker, Philip Engel, Bruno Klingler and Salim Tayou, and to thank the referee for a careful reading.
@@ 2. Reminder on linear meromorphic differential equations
on the open disc $\Delta(\rho)=\{z\in{\mathbb{C}},|z|<\rho\}$. Here $A$ is an $r\times r$-matrix with entries holomorphic functions on the punctured disc $\Delta(\rho)^{\times}=\Delta(\rho)\setminus\{0\}$ which are meromorphic along 00. In order to exclude undesirable behavior at the boundary of $\Delta(\rho)$, we assume that $A$ is overconvergent, i.e. that all its entries extend to holomorphic functions on $\Delta(\tilde{\rho})^{\times}$ for some $\tilde{\rho}>\rho$.
Recall that a gauge transformation by $P$ changes $A$ to the coefficient matrix $P^{-1}AP-zP^{-1}P^{\prime}$ in (2.1). A pullback along the finite covering $z\mapsto z^{m}$ for $m\in\mathbb{Z}_{\geq 1}$ replaces $A(z)$ by $mA(z^{m})$. We say that $A$ is in split normal form if there exist $Q_{1}(z),\ldots,Q_{\ell}(z)\in\frac{1}{z}{\mathbb{C}}[\frac{1}{z}]$ and constant square matrices $B_{1},\ldots,B_{\ell}$ of size $r_{1},\ldots,r_{\ell}$ with $\sum_{i=1}^{\ell}r_{i}=r$ such that $A$ is block diagonal with blocks $Q_{1}{\rm Id}_{r_{1}}+B_{1},\ldots,Q_{\ell}{\rm Id}_{r_{\ell}}+B_{\ell}$. Note that for such a split normal form $A$ a fundamental matrix solution on a sector has the shape
@@
  • $f$ is the restriction to $D_{d}(\rho)$ of a real analytic function defined on $D_{d}(\rho^{\prime})$ for some $\rho^{\prime}>\rho$;
@@
    for all $n\geq 0$ and all $z$ in the above sector;
  • the Taylor coefficients $\lim_{z\to 0}f_{j}^{(m)}(z)$ exist for the above $f_{j}$ for all $m\geq 0$.
@@
  • The condition on the slopes in Theorem 2.3 is satisfied after pullback along a finite covering $z\mapsto z^{m}$ for large $m$.
@@
is a fundamental matrix solution of the differential equation (2.1) on $D_{d}(\rho)$.
@@
Note here there is a change of notation as compared to (2.2), $\ell$ becomes $L$. This is because we request the $B_{i}$ to be Jordan blocks, which is achieved after a constant gauge transformation starting from the $B_{i}$ and $Q_{i}(z)$ as in (2.2). So in fact $L\geq\ell$.
@@ 3. Reminder on o-minimal structures
Recall that an o-minimal structure $(\mathcal{A}_{n})_{n}$ consists of a set of subsets $A_{n}$ of $\mathbb{R}^{n}$ which satisfy certain properties, in particular they comprise the semi-algebraic sets and $\mathcal{A}_{1}$ consists of the finite unions of intervals. In this note if not explicitly mentioned otherwise we use the o-minimal structure $\mathbb{R}_{\rm an,exp}$ which is the structure generated by the semi-algebraic sets, the graphs of real analytic functions on $[0,1]^{n}$ and the graph of $\exp\colon\mathbb{R}\to\mathbb{R}$. The key property of o-minimal structures that we use is that a real analytic definable function $f\colon(0,1)\to\mathbb{R}$ which is not identically zero has only a finite number of zeros. A basic reference for o-minimal structures is [vD98].
@@
By an overconvergent flat bundle $(\mathcal{E},\nabla)$ on $U(\rho)=(\Delta(\rho)^{\times})^{n}\times\Delta(\rho)^{d-n}$ meromorphic along $D(\rho)=\Delta(\rho)^{d}\setminus U(\rho)$ we mean the restriction to $U(\rho)$ of a flat bundle on $U(\tilde{\rho})$ for some $\tilde{\rho}>\rho$ with a meromorphic structure along $D(\tilde{\rho})$ in the sense of [Del70, II.2.13], i.e. it is provided with a prolongation to a coherent sheaf of $\mathcal{M}_{U(\tilde{\rho}),D(\tilde{\rho})}$-modules with flat connection, where $\mathcal{M}_{U(\tilde{\rho}),D(\tilde{\rho})}$ is the sheaf of rings of holomorphic functions on $U(\tilde{\rho})$ which are meromorphic along $D(\tilde{\rho})$. Such a $\mathcal{E}$ has a canonical definable structure ([BM23, Intro.]).
@@
The elementary technical observation about oscillating functions that we need in the proof of our main theorem is the following. For a complex number $z$ we write $z={\rm Re}(z)+i{\rm Im}(z)$ with ${\rm Re}(z),{\rm Im}(z)\in{\mathbb{R}}$.
@@
and grows faster than $|{\rm log}(z)|$. Thus
@@ 4. Reduction to sectors
We only have to show $\Leftarrow$”. Let $U\subset X$ be an arbitrary definable open subset and $\sigma\colon U\to\mathcal{E}|_{U}$ a flat section. Let $\overline{U}\subset\overline{X}$ be the closure of $U$. Then there are finitely many charts $\phi$ as above covering $\overline{U}$ and finitely many polysectors $S$ in these charts which cover $U$. By definability of $U$, for any such polysector $S$, the number of connected components of $U\cap S$ is finite [vD98, Sec. 2.2]. Let $S^{\circ}$ be such a connected component. We can extend the flat section $\sigma|_{S^{\circ}}$ to a flat section of $\mathcal{E}|_{S}$ which is definable by assumption. So also $\sigma|_{S^{\circ}}$ is definable. The definability of $\sigma$ on each component $S^{\circ}$ of $U\cap S$ implies the definability of $\sigma|_{U\cap S}$. ∎
@@
$\Leftarrow$” (Bakker–Mullane) By assumption the eigenvalues of $M_{1},\ldots,M_{n}$ are unitary. Let $B_{1},\ldots,B_{n}$ be commuting complex matrices with real eigenvalues and $M_{j}=e^{2\pi iB_{j}}$. Then the entries of
@@
is meromorphic along $D$, i.e. $f_{j}\equiv 0$ for $j\ll 0$. For this define $E_{j}=\{f_{j}=0\}\subset\Delta(\rho)^{d-1}$ if $f_{i}\not\equiv 0$ and $E_{j}=\varnothing$ else. This is a meager subset of $\Delta(\rho)^{d-1}$. So the countable union $\cup_{j}E_{j}$ is a meager subset of $\Delta(\rho)^{d-1}$ as well. Thus there exists $(\tilde{z}_{2},\ldots,\tilde{z}_{d})\in\Delta(\rho)^{d-1}\setminus E$. Then if $\mathcal{\tilde{Y}}(z)$ is not meromorphic along $D$, the one variable function $z_{1}\mapsto\mathcal{\tilde{Y}}(z_{1},\tilde{z}_{2},\ldots,\tilde{z}_{d})$ is not meromorphic either. This reduces us to the case $d=1$ for checking regularity. The monodromy $M_{1}$ is also calculated by restriction to these one-dimensional discs. So the implication follows from Proposition 5.1
@@ 5. Local one variable case
Our key ingredient in analyzing the local one variable case is the theory of irregular singular meromorphic differential equations and the o-minimal expansion $\mathbb{R}_{\rm sum,exp}$ of $\mathbb{R}_{\rm an,exp}$ involving the multisummable functions due to [vDS00].
@@
The coherent sheaf of $\mathcal{M}_{(\Delta(\rho)^{\times},0)}$-modules $\mathcal{E}$ on $\Delta(\rho)$ is locally free around 00, so without loss of generality $\mathcal{E}$ is the free sheaf ($\mathcal{M}^{r}_{(\Delta(\rho),0)}$. Then $(\mathcal{E},\nabla)$ corresponds to an overconvergent meromorphic differential equation (2.1), see [Del70, Sec. I.3].
Assume Proposition 5.1 is false, and say the connection is irregular singular. We can take the pullback along $z\mapsto z^{m}$ for large $m$ and perform a meromorphic gauge transformation so that we are in the situation of Corollary 2.5. Assume without loss of generality that the $Q_{1}(z)$ in this corollary is non-zero, say
@@
we obtain the $\mathbb{R}_{\rm sum,exp}$-definable function